OpenStudy (anonymous):

Compute the flux of the vector field F through the surface S. F = 2r and S is the part of the surface z = x^2 + y^2 above the disk x^2 + y^2 ≤ 4, oriented downward.

OpenStudy (anonymous):

I keep getting pi but its wrong

ganeshie8 (ganeshie8):

show ur work

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\int\limits(x i+yj+(x^2+y^2)k * (2 x i +2yj - k)dxdy\]

OpenStudy (anonymous):

\[\int\limits (2x^2 +2y^2 -x^2 -y^2)dxdy\]

OpenStudy (anonymous):

\[\int\limits(x^2+y^2) dxdy\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi} \int\limits_{0}^{2} r^2rdrd \theta\]

OpenStudy (anonymous):

I got 8pi this time but its wrong

OpenStudy (anonymous):

@ganeshie8 what do you think i'm doing wrong

ganeshie8 (ganeshie8):

try -8pi

ganeshie8 (ganeshie8):

\[\int\langle x i+yj+(x^2+y^2\rangle \cdot \langle -2 x i -2yj+ k \rangle ~dxdy \]

ganeshie8 (ganeshie8):

cuz the surface is oriented downward

ganeshie8 (ganeshie8):

\[\iint\langle x i+yj+(x^2+y^2\rangle \cdot \langle -2 x i -2yj+ k \rangle ~dxdy \] btw its good to use double integral whenever you mean a double integral

OpenStudy (anonymous):

well that why is positive

ganeshie8 (ganeshie8):

-8pi didnt work ?

OpenStudy (anonymous):

this is the formula

ganeshie8 (ganeshie8):

it wont open for me

OpenStudy (anonymous):

the formula youre using is for oriented upward

ganeshie8 (ganeshie8):

let me lookup my notes

ganeshie8 (ganeshie8):

have u tried -8pi ?

OpenStudy (anonymous):

\[\int\limits \int\limits F(x,y, f(x,y) ) * (f_x i +f_yj -k)dxdy \]

OpenStudy (anonymous):

this is the formula for oriented downward

ganeshie8 (ganeshie8):

nvm, you're right upward : <-fx, -fy, 1> downward <fx, fy, -1>

OpenStudy (anonymous):

I'm wondering if my integral points are right

ganeshie8 (ganeshie8):

hey F = 2r right?

ganeshie8 (ganeshie8):

you're using F=r instead it seems

OpenStudy (anonymous):

yeah F=2r

OpenStudy (anonymous):

that's why I'm going from 0 to 2

ganeshie8 (ganeshie8):

what has your region got to do anything with vectorfield ?

OpenStudy (anonymous):

nothing, i'm confuse

ganeshie8 (ganeshie8):

try 16pi

OpenStudy (anonymous):

what does the

OpenStudy (anonymous):

I only have one more try

ganeshie8 (ganeshie8):

your bounds depend only on the shadow in xy plane

ganeshie8 (ganeshie8):

0->2 0->2pi give you the full disk of radius 2 in xy plane

ganeshie8 (ganeshie8):

they are not related to given vector field

ganeshie8 (ganeshie8):

F = 2r = <2x, 2y, 2z>

OpenStudy (anonymous):

so it's\[\int\limits_{0}^{2\pi} \int\limits_{0}^{1}\]

ganeshie8 (ganeshie8):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^2\langle 2x i+2yj+2(x^2+y^2\rangle \cdot \langle 2 x i +2yj- k \rangle ~dxdy \]

OpenStudy (anonymous):

so why is it \[\int\limits_{0}^{}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{?}\]

OpenStudy (anonymous):

0 to 2

OpenStudy (anonymous):

\[\int\limits_{0}^{2}\]

OpenStudy (anonymous):

becuae 0f <4?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

how do u represent the disk x^2+y^2 <= 4 in polar ?

OpenStudy (anonymous):

so the radius is 2

OpenStudy (anonymous):

so what about if \[x^2 +x^2 \le1\]

OpenStudy (anonymous):

that would make it \[\int\limits_{0}^{1}\]

ganeshie8 (ganeshie8):

thats right for representing the radius "r", it ranges from 0 to 1 if you are on disk of radius 1

OpenStudy (anonymous):

so what's making it 0 to 2 in this example? why not 0 to 4?

ganeshie8 (ganeshie8):

whats the radius of circle x^2+y^2 = 4 ?

OpenStudy (anonymous):

oh yeah 2 haha the diameter is 4

ganeshie8 (ganeshie8):

right

OpenStudy (anonymous):

so \[\int\limits \int\limits 2(x^2+y^2)dxdy\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2} 2r^2rdr d \theta\]

ganeshie8 (ganeshie8):

looks good to me